a) find the kumet takes bcfore the stone Anather Stone as thrn uertica
Answer #1

Answer:

(a) Given that t = 0 s and initial speed u = 20 m/s

using the kinematics equation v = u + at,

In the upward motion a = -g then v = u - gt, At maximum height the speed of the stone becomes zero, that means final speed is zero i.e., v = 0

then 0 = u - gt or u = gt   or t = u / g and let us assume this as t1 = u/g

and using the expression s = ut + 1/2gt2, for free fall body u = 0 then s = 1/2gt2 and using the value t = u/g then

s = 1/2 g(u/g)2 = u2/2g = maximum height H,

After reaching maximum height, the stone begins to travel downward like a free fall object. So height h = 1/2 gt22 or

t22 = 2h/g = 2 (u2/2g)/g = u2/g2, therefore t2 = u/g

From this we can say that the time of ascent is equal to the time of descent

Therefore, total time t = t1 + t2 = u/g + u/g = 2u/g

Then t = 2 (20 m/s) / (9.8 m/s2) = 4.08 s

(b) Maximum height reached by the stone is hmax = u2/2g = (20 m/s)2/ 2 (9.8 m/s2) = 20.40 m.

(c) Two stones thrown vertically upwards with the same speed 20 m/s but with the time difference 2 s

then using the expressions , s1 = ut - 1/2 gt2 and s2 = u(t-2) - 1/2 g(t-2)2

Equating these two equations then we obtain

ut - 1/2 gt2 = u(t-2) - 1/2 g(t-2)2

2u = 1/2 gt2 - 1/2 g(t-2)2 = g/2 [ t2 - (t-2)2 ] = g/2 [ t2 - (t2 + 4 - 4t) ] = g/2 [4t - 4] = g (2t - 2)

Therefore, 2u = g (2t - 2) or u = g/2 (2t - 2)

Substituting the values then we get,

20 m/s = (9.8 m/s2)/2 (2t - 2)

4.08 s = 2t - 2 or t = 3.04 s at this time the two stones will meet.

using the equation, s1 = ut - 1/2 gt2

height s1 = (20 m/s) (3.04 s) - 1/2 (9.8 m/s2) (3.04 s)2 = 60.81 m - 45.28 m = 16.53 m at this point the two stones will meet.